OpenStudy (anonymous):

Below are the data collected from two random samples of 100 members of a large travel club regarding the type of vacation they prefer: Sample Adventure Beach Cruise Ski A 73 4 18 3 B 71 6 22 1

OpenStudy (anonymous):

@wolf1728 @Michele_Laino @Abhisar

OpenStudy (anonymous):

Which of the following inferences can be made based on the data? Most members prefer a beach vacation. Most members prefer a cruise vacation. More members prefer a cruise vacation and a ski vacation than a beach vacation. More members prefer a beach vacation and a cruise vacation than an adventure vacation.

OpenStudy (anonymous):

@Michele_Laino i think its C. not sure

OpenStudy (michele_laino):

you have to compute the singles frequencies, namely the frequencies for a beach vacation are: 4/100 =0.04 for group A 6/100=0.06 for group B by the way group A contains 98 elements

OpenStudy (anonymous):

OK

OpenStudy (michele_laino):

fro cruise vacation we have: 18/100=0.18 for group A 22/100=0.22 for group B

OpenStudy (michele_laino):

for cruise vacation and ski vacation we have: (18+3)/100=21/100=0.21 group A (22+1)/100= 23/100=0.23 group B

OpenStudy (anonymous):

oh OK so would my answer then be D. @Michele_Laino

OpenStudy (anonymous):

well actually no cause an Adventure vacation is preferred alot :(

OpenStudy (michele_laino):

for beach and cruise vacation we have: (18+4)/100= 0.22 group A (22+6)/100= 0.28 group B and adventure vacation: 73/100=0.73 group A 71/100=0.71 group B

OpenStudy (michele_laino):

answer D is not correct, since 0.73>0.22 and 0.71>0.28, am I right?

OpenStudy (michele_laino):

whereas: 0.21 > 0.04, and 0.23> 0.06 so, what can you conclude?

OpenStudy (anonymous):

was i right then about C. @Michele_Laino

OpenStudy (michele_laino):

yes! since for beach vacation and cruise vacation the frequencies are less than 0.5

OpenStudy (anonymous):

Thankk You!!! could u please help me with a few more? @Michele_Laino

OpenStudy (michele_laino):

ok!